# Polariton assisted incoherent to coherent excitation energy transfer between colloidal nanocrystal quantum dots

### v1

@@ I Introduction
In the realm of nanophotonics and quantum optics, understanding and controlling the transfer of energy between quantum emitters is paramount for advancing technologies ranging from solar cells to quantum information processing.[1, 2, 3] Colloidal nanocrystal (NC) quantum dots (QDs), with unique electronic and optical properties,[4, 5, 6] have emerged as versatile platforms for studying energy transfer processes at the nanoscale.[7, 1, 8, 9] Among these processes, Fluorescence Resonance Energy Transfer (FRET) has garnered considerable attention for its role in mediating efficient energy transfer between quantum dots[1, 10, 9] and other emitters[11, 12] and has been extensively studied and exploited in various fields, including molecular biology,[13, 14] where it serves as a fundamental tool for probing molecular interactions and dynamics.[12, 15, 16] FRET, an incoherent energy transfer process,[17, 18] relies on the weak coupling between a donor and an acceptor fluorophore, with a transfer rate given by Fermi’s golden rule (FGR):[19, 20]
@@
Here, $\rho_{{\cal D}}$ is the population distribution of excited states on the donor, $J_{{\cal DA}}$ is the electromagnetic coupling between transition densities connecting the ground state and excited state for the donor and the acceptor (approximated by the dipole-dipole term, see Eq. (3) below), and $E_{{\cal D}}$ and $E_{{\cal A}}$ are the donor and acceptor excited state energies, respectively. When the emission spectrum of the donor overlaps with the absorption spectrum of the acceptor and they are within a certain distance, typically within a few nanometers, energy transfer can occur with high efficiency, but on relatively long timescales (a few tens of nanoseconds).[21, 11, 20, 10, 22, 23] FRET processes are inherently limited by the steep distance dependence ($R^{-3}$) of the dipole-dipole coupling between the donor and acceptor QDs, resulting in small couplings, overdamped energy transfer dynamics, and long transfer rates, necessitating the exploration of alternative mechanisms to accelerate and control the energy transfer mechanism.[24, 25, 26]
One strategy for regulating the exciton energy transfer process involves utilizing a donor-bridge-acceptor configuration, wherein the transition mechanism and rates can be adjusted by manipulating the states of the bridge.[27, 28] Recent advancements have led to the emergence of polariton–mediated energy transfer as a promising alternative to FRET.[29, 25, 30] Polaritons, arising from the strong coupling between photons and excitons, present distinct advantages for facilitating energy transfer, such as heightened coupling strength and diminished sensitivity to spectral overlap.[25, 30, 31, 26, 32] As efforts intensify to exploit polariton-mediated energy transfer, ongoing research endeavors are dedicated to elucidating the control parameters dictating the efficiency and dynamics of these processes.[33, 24, 26, 34]
@@ II.1 Donor-acceptor and cavity Hamiltonians
$+\sum_{\alpha\in\mathcal{D},\mathcal{A}}\hbar\omega_{\alpha}b_{\alpha}^{\dagger}b_{\alpha}+\sum_{n,m\in\mathcal{D},\mathcal{A}}V_{nm}^{\alpha}\left|\psi_{n}\right\rangle\left\langle\psi_{m}\right|q_{\alpha}.$ (2)
The first and second terms on the right hand side describe the manifold of donor and acceptor excitonic states and their couplings, respectively, and the remaining terms describe the lattice vibrations and the exciton–phonon couplings, approximated to lowest order in the phonons displacement, $q_{\alpha}$. In the above, $E_{n}$ is the energy of a singly excited state $\left|\psi_{n}\right\rangle$, with $n\in\mathcal{D}/\mathcal{A}$ and $\mathcal{D}=\left\{{\cal D}_{1},{\cal D}_{2},\cdots\right\}$, $\mathcal{A}=\left\{{\cal A}_{1},{\cal A}_{2},\cdots\right\}$. We use the notation $\left|\psi_{{\cal D}_{1}}\right\rangle$ to describe the donor in excitonic state ${\cal D}_{1}$ while the acceptor is in its ground state and vise versa for $\left|\psi_{{\cal A}_{1}}\right\rangle$. The excitonic energies and wave functions ($E_{n}$ and $\left|\psi_{n}\right\rangle$) were obtained by ignoring the coupling between the donor and acceptor, using the semi–empirical pseudopotential methods combined with the Bethe–Salpeter equation (BSE).[37] The quasiparticle states near the top of the valence band and the bottom of the conduction band were generated for each QD using the filter diagonalization technique[39, 40] with a real–space grid basis of $N_{g}\approx 2\cdot 10^{6}$ and a grid spacing of $\approx 0.75a_{0}$. A total of $80$ electron (unoccupied) and $140$ hole (occupied) states were generated to construct the Bethe−Salpeter Hamiltonian,[41] within the static screening approximation,[37] providing converged excitonic energies and transition dipole moments for the lowest $30$ excitonic states that participate in the dynamics, for each QD.
@@
$J_{\mathcal{DA}}=f_{{\cal D}}f_{{\cal A}}\frac{\left|\boldsymbol{\mu}_{{\cal D}}^{{\rm ex}}\right|\left|\boldsymbol{\mu}_{{\cal A}}^{{\rm ex}}\right|}{\epsilon_{\text{m}}\left|R\right|^{3}}\kappa\left(\theta_{{\cal DA}},\theta_{{\cal D}},\theta_{{\cal A}}\right),$ (3)
@@
The donor and acceptor excitonic states also couple to the lattice vibrations with frequencies $\omega_{\alpha}$ and modes $q_{\alpha}=\sqrt{\frac{\hbar}{2\omega_{\alpha}}}\left(b_{\alpha}^{\text{$\dagger$}}+b_{\alpha}\right)$, obtained by diagonalizing the dynamical matrix of the QD donor and acceptor using the Stilling–Weber force field parameterized for Cd, Se, S elements.[45] The exciton–phonon couplings were expanded to lowest order in the displacements, $q_{\alpha}$, and the strength of coupling between exciton $\left|\psi_{\mathcal{D}/{\cal A}}\right\rangle$ to mode $\alpha$ was obtained from the pseudopotential Hamiltonian.[36] We assumed that the excitons are localized to the donor or the acceptor and thus, excitons only couple to the lattice modes of the donor or the acceptor. The coupling of excitons to the lattice vibrations is key in accounting for the reorganization and the relaxation of the system to thermal equilibrium.
@@
where $a$ ($a^{\dagger}$) is the annihilation (creation) operator of a photon with frequency $\omega_{c}$, and $\hbar g_{n}$ is coupling strength between the cavity mode and an excitonic state $\left|\psi_{n}\right\rangle$ ($n\in\mathcal{D}/\mathcal{A}$). The cavity–exciton coupling strength is given by $g_{n}=\sqrt{\omega_{c}/\left(2\hbar\epsilon_{\text{m}}\mathcal{V}\right)}\boldsymbol{\mu}_{n}^{{\rm ex}}\cdot\boldsymbol{k}$, where as before, $\boldsymbol{\mu}_{n}^{{\rm ex}}$ is the transition dipole moment from the ground state $\left|\psi_{g}\right\rangle$ to exciton $\left|\psi_{n}\right\rangle$ (obtained from the BSE calculation) and $\boldsymbol{k}$ is a unit vector of the polarization direction of the electromagnetic field. $\epsilon_{{\rm m}}$ is the effective permittivity inside the cavity (medium) and $\mathcal{V}$ is the effective cavity quantization volume.
@@ II.2 Polaritonic representation and polaron transformation
The total Hamiltonian can also be expressed using the polaritonic bases, $\left\{\left|\varphi_{n}\right\rangle=c_{ng}\left|\psi_{g},1\right\rangle+\sum_{m}c_{nm}^{\text{ex}}\left|\psi_{m},0\right\rangle\right\}$, where $\left|\psi_{g},1\right\rangle$ and $\left|\psi_{m},0\right\rangle$ are the ground state with one photon or an excitonic state with zero photons, respectively. This representation simplifies the description of the dynamics within the Redfield approach described below, and is not limited by the weak donor-acceptor coupling strength (provided that the exciton-phonon couplings are small), as is the case for FRET between CdSe/CdS core-shell QDs. The coefficients $c_{ng}$ and $c_{nm}^{\text{ex}}$ were obtained by diagonalizing the system Hamiltonian, defined by (we assume that the ground state energy, $E_{g}=0$):
@@
In addition, to properly describe dephasing and relaxation and to account for ultrafast multi–phonon relaxation channels, we performed a small polaron transformation as detailed in Ref. 38, 47. The total transformed Hamiltonian ${\cal H}$ in the polaritonic basis is given
@@
In the above, the dressed coupling is given by
@@
As evident, the dressed coupling depends exponentially on the momenta of all phonon modes, enabling multi–phonon relaxation, even at the lowest order perturbation in $\mathcal{H}_{\text{I}}$. In the above equations, $\tilde{E}_{n}$ and $\lambda_{n}=\frac{1}{2}\sum_{\alpha}\left(\tilde{V}_{nn}^{\alpha}\right)^{2}/\omega_{\alpha}^{2}$ are the energy and reorganization energy (polaron shift) of polaritonic state $\left|\varphi_{n}\right\rangle$, respectively. The reorganization energy depends on $\tilde{V}_{nm}^{\alpha}$, the coupling matrix element between two polaritonic states $\left|\varphi_{n}\right\rangle$ and $\left|\varphi_{m}\right\rangle$ and phonon mode $\alpha$. $\tilde{V}_{nm}^{\alpha}$ was obtained by applying the unitary transformation $U$ that diagonalizes $H_{{\rm S}}$ to $\tilde{V}^{\alpha}=U^{\dagger}V^{\alpha}U$, for each mode $\alpha$. $p_{\alpha}$ is the momentum operator of vibrational mode $\alpha$.
@@ II.3 Dynamics
-- $\left.C_{nm,kl}^{*}\left(\tau\right)\left[\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{nm},e^{-i/\hbar\mathcal{H}_{\text{S}}\tau}\sigma\left(t-\tau\right)\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{kl}e^{i/\hbar\mathcal{H}_{\text{S}}\tau}\right]\right\},$ (8)
where the polariton-phonon coupling is expressed as $\mathcal{H}_{\text{I}}=\sum_{nm}W_{nm}\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{nm}$ with $\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{nm}=\left(1-\delta_{nm}\right)\left|\varphi_{n}\right\rangle\left\langle\varphi_{m}\right|$. In the above equation, $\left\langle\cdots\right\rangle_{\text{B}}$ denotes an equilibrium thermal average over the so called “bath” degrees of freedom. The bath correlation function appearing above given by:
@@ III Exciton energy transfer between QDs
$\rho_{{\cal D}}\left(t\right)=\left(1-P_{{\cal D}}^{\text{eq}}\right)\exp\left[-k_{\text{ET}}t\right]+P_{{\cal D}}^{\text{eq}}$ (11)
@@ IV Polaritonic energy transfer between QDs
where $\omega_{{\rm nu}}=\sqrt{\frac{\sum_{\alpha}\omega_{\alpha}^{2}\lambda_{\alpha}}{\sum_{\alpha}\lambda_{\alpha}}}$ is the average phonon frequency weighted by the reorganization energy for each phonon mode. The terms $\lambda_{{\cal D}}$ and $\lambda_{{\cal A}}$ represent the average reorganization energies for the donor and acceptor, respectively, defined as $\lambda_{{\cal D}/{\cal A}}=\sum_{n\in{\cal D}/{\cal A}}P_{n}\lambda_{n}$, where $P_{n}$ is the Boltzmann population distribution of exciton $|\psi_{n}\rangle$ and, as before, $\lambda_{n}=\frac{1}{2}\sum_{\alpha}\left(\tilde{V}_{nn}^{\alpha}\right)^{2}/\omega_{\alpha}^{2}$. In Fig.  5(a) we plot the nonadiabatic parameter $\gamma$ as a function of the QD-cavity coupling strength $\hbar g$. For small coupling to the cavity mode, $\gamma\ll 1$ and the system is in the nonadiabatic regime, where the dynamics are characterized by an exponential decay of populations. As the coupling to the cavity mode increases, $\gamma$ also increases due to the increase in the effective coupling between the two QDs. This parameter reaches and surpasses $\gamma=1$ when $\hbar g=40$ meV, marking the transition to an adiabatic regime. In this regime, the effective coupling strength between the QDs becomes significantly larger than the couplings of the donor and acceptor polaritonic states to phonons, leading to more pronounced and sustained coherent dynamics.
In Fig. 5(b), we present the exciton transfer rate as a function of the effective coupling for both ideal and lossy cavities. For small value of $J_{\text{eff}}$, the transfer rate scales with $J_{\text{eff}}^{2}$, consistent with Fermi’s golden rule. As $J_{\text{eff}}$ increases, the system transitions to the adiabatic regime, and the transfer rate which is determined by decoherence times exhibits a more moderate dependence, fitting to a $\sqrt{J_{\text{eff}}}$ relationship. The gray dashed line in the figure marks the crossover coupling point, where the adiabatic parameter $\gamma=1$. In the lossy cavity, the energy transfer rate is inherently slower when compared to an ideal, lossless cavity. However, as the coupling to the cavity mode is increased, this discrepancy becomes less significant. This is because the rate of energy transfer accelerates, eventually becoming so rapid that it surpasses the cavity’s inherent loss rate. As a result, the effects of the cavity’s losses are effectively mitigated, rendering them negligible in the context of the swift energy transfer times that are achieved.

### v4

@@ I Introduction
In the realm of nanophotonics and quantum optics, understanding and controlling the transfer of energy between quantum emitters is paramount for advancing technologies ranging from solar cells to quantum information processing.[1, 2] Colloidal nanocrystal (NC) quantum dots (QDs), with unique electronic and optical properties,[3, 4, 5] have emerged as versatile platforms for studying energy transfer processes at the nanoscale.[6, 1, 7, 8] Among these processes, Fluorescence Resonance Energy Transfer (FRET) has garnered considerable attention for its role in mediating efficient energy transfer between quantum dots[1, 9, 8] and other emitters,[10, 11] and has been extensively studied and exploited in various fields, including molecular biology,[12, 13] where it serves as a fundamental tool for probing molecular interactions and dynamics.[11, 14] FRET, an incoherent energy transfer process,[15, 16] relies on the weak coupling between a donor and an acceptor fluorophore, with a transfer rate given by Fermi’s golden rule (FGR):[17]
@@
Here, $\rho_{{\cal D}}$ is the population distribution of excited states on the donor, $J_{{\cal DA}}$ is the electromagnetic coupling between transition densities connecting the ground state and excited state for the donor and the acceptor (approximated by the dipole-dipole term, see Eq. (3) below), and $E_{{\cal D}}$ and $E_{{\cal A}}$ are the donor and acceptor excited state energies, respectively. When the emission spectrum of the donor overlaps with the absorption spectrum of the acceptor and they are within a certain distance, typically within a few nanometers, energy transfer can occur with high efficiency, but on relatively long timescales (a few tens of nanoseconds).[10, 9, 18] FRET processes are inherently limited by the steep distance dependence ($R^{-3}$) of the dipole-dipole coupling between the donor and acceptor QDs, resulting in small couplings, overdamped energy transfer dynamics, and long transfer rates, necessitating the exploration of alternative mechanisms to accelerate and control the energy transfer mechanism.[19, 20, 21]
One strategy for regulating the exciton energy transfer process involves utilizing a donor–bridge–acceptor configuration, wherein the transition mechanism and rates can be adjusted by manipulating the states of the bridge.[22, 23] Recent advancements have led to the emergence of polariton–mediated energy transfer as a promising alternative to FRET.[24, 20, 25, 26] Polaritons, arising from the strong coupling between photons and excitons,[27] present distinct advantages for facilitating energy transfer, including heightened coupling strength, tunable spectral overlap, and minimal dependence on distance, making efficient remote energy transfer possible with more flexible material design.[28, 26, 29, 21, 30] As efforts intensify to exploit polariton-mediated energy transfer, ongoing research endeavors are dedicated to elucidating the control parameters dictating the efficiency and dynamics of these processes, such as energy decay rate, dephasing rate, and Rabi splitting/light coupling strength, as well as their scaling with number of molecules involved.[31, 32, 21, 33]
@@ II.1 Donor-acceptor and cavity Hamiltonians
$+\sum_{\alpha}\hbar\omega_{\alpha}b_{\alpha}^{\dagger}b_{\alpha}+\sum_{\alpha}\sum_{n,m\in\mathcal{D},\mathcal{A}}V_{nm}^{\alpha}\left|\psi_{n}\right\rangle\left\langle\psi_{m}\right|q_{\alpha}.$ (2)
The first and second terms on the right hand side describe the manifold of donor and acceptor excitonic states and their couplings, respectively, and the remaining terms describe the lattice vibrations and the exciton–phonon couplings, approximated to lowest order in the phonons displacement, $q_{\alpha}$. In the above, $E_{n}$ is the energy of a singly excited state $\left|\psi_{n}\right\rangle$, with $n\in\mathcal{D}/\mathcal{A}$ and $\mathcal{D}=\left\{{\cal D}_{1},{\cal D}_{2},\cdots\right\}$, $\mathcal{A}=\left\{{\cal A}_{1},{\cal A}_{2},\cdots\right\}$. We use the notation $\left|\psi_{{\cal D}_{1}}\right\rangle$ to describe the donor in excitonic state ${\cal D}_{1}$ while the acceptor is in its ground state and vise versa for $\left|\psi_{{\cal A}_{1}}\right\rangle$. The excitonic energies and wave functions ($E_{n}$ and $\left|\psi_{n}\right\rangle$) were obtained by ignoring the coupling between the donor and acceptor, using the semi–empirical pseudopotential methods combined with the Bethe–Salpeter equation (BSE).[36] The quasiparticle states near the top of the valence band and the bottom of the conduction band were generated for each QD using the filter diagonalization technique[38, 39] with a real–space grid basis of $N_{g}\approx 2\cdot 10^{6}$ and a grid spacing of $\approx 0.75a_{0}$. A total of $80$ electron (unoccupied) and $140$ hole (occupied) states were generated to construct the Bethe–Salpeter Hamiltonian,[40] within the static screening approximation,[36] providing converged excitonic energies and transition dipole moments for the lowest $30$ excitonic states that participate in the dynamics, for each QD.
@@
$J_{\mathcal{DA}}=f_{{\cal D}}f_{{\cal A}}\frac{\left|\boldsymbol{\mu}_{{\cal D}}^{{\rm ex}}\right|\left|\boldsymbol{\mu}_{{\cal A}}^{{\rm ex}}\right|}{\epsilon_{\text{m}}\left|R\right|^{3}}\kappa\left(\theta_{{\cal DA}},\theta_{{\cal D}},\theta_{{\cal A}}\right)$ (3)
@@
The donor and acceptor excitonic states also couple to the lattice vibrations with frequencies $\omega_{\alpha}$ and modes $q_{\alpha}=\sqrt{\frac{\hbar}{2\omega_{\alpha}}}\left(b_{\alpha}^{\text{$\dagger$}}+b_{\alpha}\right)$, obtained by diagonalizing the dynamical matrix of the QD donor and acceptor using the Stilling–Weber force field parameterized for Cd, Se, S elements.[44] The exciton–phonon couplings were expanded to lowest order in the displacements, $q_{\alpha}$ and the strength of coupling between exciton $\left|\psi_{n}\right\rangle$ and $\left|\psi_{m}\right\rangle$ to mode $\alpha$, $V_{nm}^{\alpha}$, was obtained from the pseudopotential Hamiltonian.[35] We considered both diagonal $V_{n=m}^{\alpha}$ and off-diagonal $V_{n\neq m}^{\alpha}$ coupling terms and assumed that the excitons couple to the lattice modes on either the donor or the acceptor, depending on whether they localize to the donor or the acceptor. The coupling of excitons to the lattice vibrations is key in accounting for the reorganization and the relaxation of the system to thermal equilibrium.
@@
where $a$ ($a^{\dagger}$) is the annihilation (creation) operator of a photon with frequency $\omega_{c}$, and $\hbar g_{n}$ is coupling strength between the cavity mode and an excitonic state $\left|\psi_{n}\right\rangle$ ($n\in\mathcal{D}/\mathcal{A}$). The cavity–exciton coupling strength is given by $g_{n}=\sqrt{\omega_{c}/\left(2\hbar\epsilon_{\text{m}}\mathcal{V}\right)}\boldsymbol{\mu}_{n}^{{\rm ex}}\cdot\boldsymbol{\varepsilon}$, where as before, $\boldsymbol{\mu}_{n}^{{\rm ex}}$ is the transition dipole moment from the ground state $\left|\psi_{g}\right\rangle$ to exciton $\left|\psi_{n}\right\rangle$ (obtained from the BSE calculation) and $\boldsymbol{\varepsilon}$ is a unit vector of the polarization direction of the electromagnetic field. $\epsilon_{{\rm m}}$ is the effective permittivity inside the cavity (medium) and $\mathcal{V}$ is the effective cavity quantization volume. The last two terms in the above equation correspond to the quadratic dipole self-energy, which has shown to be important in the ultra-strong coupling limit.[47] The dipole self-energy term arises from the Power-Zienau-Woolley transformation that explicitly mixes the electronic and photonic degrees of freedom.[48] For the results reported in this work, the effect of the dipole-self-energy is negligible since $\omega_{c}\gg g_{n}$.
@@ II.2 Polaritonic representation and polaron transformation
The total Hamiltonian can also be expressed using the polaritonic basis, $\left\{\left|\varphi_{n}\right\rangle=c_{ng}\left|\psi_{g},1\right\rangle+\sum_{m}c_{nm}^{\text{ex}}\left|\psi_{m},0\right\rangle\right\}$, where $\left|\psi_{g},1\right\rangle$ and $\left|\psi_{m},0\right\rangle$ are the ground state with one photon or an excitonic state with zero photons, respectively. This representation simplifies the description of the dynamics within the Redfield approach described below, and is not limited by the weak donor-acceptor coupling strength (provided that the exciton-phonon couplings are small), as is the case for FRET between CdSe/CdS core-shell QDs. The coefficients $c_{ng}$ and $c_{nm}^{\text{ex}}$ were obtained by diagonalizing the system Hamiltonian (describing only the excitons and photons without the phonons), defined by:
@@
In the above, we assumed that the ground state energy is zero, namely $E_{g}=0$. Using the polaritonic basis obtained by diagonalizing the system Hamiltonian in Eq. (5), the total Hamiltonian ($\mathcal{H}=\mathcal{H}_{{\rm S}}+\mathcal{H}_{\text{B}}+\mathcal{H}_{\text{I}}$) given by $H_{{\rm\mathcal{DA}}}+H_{{\rm cav}}$ can be written as:
@@
In the above, $\tilde{E}_{n}$ is the polaritonic energy of state $\left|\varphi_{n}\right\rangle$ (an eigenvalue of $H_{{\rm S}}$ in Eq. (5)) and $\lambda_{n}=\frac{1}{2}\sum_{\alpha}\left(\tilde{V}_{nn}^{\alpha}\right)^{2}/\omega_{\alpha}^{2}$ is the corresponding reorganization energy (polaron shift). The dressed coupling between polaritons and phonons, $W_{nm}$, is given by:
@@
To derive the Hamiltonian in Eq. (6) we also performed a small polaron transformation (in addition to diagonalizing the system Hamiltonian to form the polaritonic states), as detailed in Refs. 37 and 46. As a result of the polaron transformation, the dressed coupling between polaritons, $W_{nm}$, depends exponentially on the momenta ($p_{\alpha}$) and position ($q_{\alpha}$) of all phonon modes, enabling the description of multi–phonon relaxation channels within the lowest order perturbation in $\mathcal{H}_{\text{I}}$. In the above, the dressed coupling (and the reorganization energy) depend on $\tilde{V}_{nm}^{\alpha}$, the transformed coupling between two polaritonic states $\left|\varphi_{n}\right\rangle$ and $\left|\varphi_{m}\right\rangle$ and phonon mode $\alpha$, obtained by applying the unitary transformation $U$ that diagonalizes $H_{{\rm S}}$ to $\tilde{V}^{\alpha}=U^{\dagger}V^{\alpha}U$, for each mode $\alpha$.
@@ II.3 Dynamics
-- $\left.C_{nm,kl}^{*}\left(\tau\right)\left[\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{nm},e^{-i/\hbar\mathcal{H}_{\text{S}}\tau}\sigma\left(t-\tau\right)\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{kl}e^{i/\hbar\mathcal{H}_{\text{S}}\tau}\right]\right\}+\mathcal{L_{\rm loss}}\sigma(t),$ (8)
where the polariton-phonon coupling is expressed as $\mathcal{H}_{\text{I}}=\sum_{nm}W_{nm}\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{nm}$ with $\left(\mathcal{H}_{\text{S}}^{{\rm I}}\right)_{nm}=\left(1-\delta_{nm}\right)\left|\varphi_{n}\right\rangle\left\langle\varphi_{m}\right|$. In the above equation, $\left\langle\cdots\right\rangle_{\text{B}}$ denotes an equilibrium thermal average over the so called ”bath” degrees of freedom. The bath correlation function appearing above is given by:
@@
In addition, to describe cavity losses, we added a phenomenological loss term represented by the last term on the right hand side of Eq. (8), given by:
$\left[\mathcal{L_{\text{loss}}}\sigma\left(t\right)\right]_{nm}=-\delta_{nm}\left(\frac{c_{ng}^{2}}{\tau^{\text{cav}}}\right)\sigma_{nm}\left(t\right),$ (10)
where $\sigma_{nm}\left(t\right)=\left\langle\varphi_{n}\left|\sigma\left(t\right)\right|\varphi_{m}\right\rangle$, and the cavity loss rate from polaritonic state $\left|\varphi_{n}\right\rangle$ is give by the ratio of the cavity fraction of the polaritonic state, $c_{ng}^{2}=\left|\left\langle\psi_{g},1|\varphi_{n}\right\rangle\right|^{2}$, and the cavity lifetime, $\tau^{\text{cav}}$.
@@ III Exciton energy transfer between QDs
$\rho_{{\cal D}}\left(t\right)=\left(1-P_{{\cal D}}^{\text{eq}}\right)\exp\left[-k_{\text{ET}}t\right]+P_{{\cal D}}^{\text{eq}},$ (12)
@@ IV Polaritonic energy transfer between QDs
where $\omega_{{\rm nu}}$ is the averaged effective phonon modes that contribute most to the system dynamics. The terms $\lambda_{{\cal D}}$ and $\lambda_{{\cal A}}$ represent the average reorganization energies for the donor and acceptor, respectively, defined as $\lambda_{{\cal D}/{\cal A}}=\sum_{n\in{\cal D}/{\cal A}}P_{n}\lambda_{n}$, where $P_{n}$ is the Boltzmann population distribution of exciton $|\psi_{n}\rangle$ and, as before, $\lambda_{n}=\frac{1}{2}\sum_{\alpha}\left(\tilde{V}_{nn}^{\alpha}\right)^{2}/\omega_{\alpha}^{2}$. In Fig. 5(a) we plot the nonadiabatic parameter $\gamma$ as a function of the QD-cavity coupling strength $\hbar g$. For small coupling to the cavity mode, $\gamma\ll 1$ and the system is in the nonadiabatic regime, where the dynamics are characterized by an exponential decay of populations. As the coupling to the cavity mode increases, $\gamma$ also increases due to the increase in the effective coupling between the two QDs. This parameter reaches and surpasses $\gamma=1$ when $\hbar g=20$ meV, marking the transition to an adiabatic regime. In this regime, the effective coupling strength between the QDs becomes significantly larger than the couplings of the donor and acceptor polaritonic states to phonons, leading to more pronounced and sustained coherent dynamics.
In Fig. 5(b), we present the exciton transfer rate as a function of the effective coupling for both ideal and lossy cavities. For small value of $J_{\text{eff}}$, the transfer rate scales with $J_{\text{eff}}^{2}$, consistent with Fermi’s golden rule. As $J_{\text{eff}}$ increases, the system transitions to the adiabatic regime, and the transfer rate which is determined by decoherence times exhibits a more moderate dependence, fitting to a $\sqrt{J_{\text{eff}}}$ relationship. In this limit, the transfer rate was obtained by fitting the envelop function of $\rho_{{\cal D}}\left(t\right)$/$\rho_{{\cal A}}\left(t\right)$ to an exponential decay/rise. The gray dashed line in the figure marks the crossover coupling point, where the adiabatic parameter $\gamma=1$. In the lossy cavity, the energy transfer rate is inherently slower when compared to an ideal, lossless cavity. However, as the coupling to the cavity mode is increased, this discrepancy becomes less significant. This is because the rate of energy transfer accelerates, eventually becoming so rapid that it surpasses the cavity’s inherent loss rate. As a result, the effects of the cavity’s losses are effectively mitigated, rendering them negligible in the context of the swift energy transfer times that are achieved.
Before we conclude, we would like to mention that our calculations have been performed for a single QD pair and such large couplings required to drive the system into the adiabatic limit are typically achieved in plasmonic nanocavities, with complications arising from the coupling of excitons to near-fields.[66] While our approach ignores the role of near fields, it provides the guidelines that would be useful in future studies of QDs in nanocavities.